# Cultural Transmission, Technology, and Treatment of the Elderly

### v1

@@ Appendix A Proof of Proposition 4
We describe the equilibrium solution to the maximization problem (35). Our general plan of attack is to proceed as if the solution to the problem is interior and then to find this solution. The solution then allows us to check when it is in fact interior. We start by assuming that agents maintain the belief that in equilibrium gifts are some multiplicative function of the value of $\eta_{t}$ and model parameters as follows:
$g_{mt+1}=\eta_{t+1}\Psi_{t+1}$ (A.1)
@@
where $\Psi_{t+1}$ is to be determined. Substituting for $g_{mt+1}$ using (A.1) into the middle-aged agent’s optimization problem (35), we find that the first-order conditions associated with an interior solution are:
$-\frac{1-\eta_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}}+\frac{\eta_{t}}{g_{mt}}$ $=0$ (A.2)
@@
The first expression in (A.2) implies that gifts are proportional to income, after inculcative expenditures and savings are subtracted out:
$g_{mt}=\eta_{t}(y_{mt}-d_{t}\eta_{t+1}-s_{mt})$ (A.3)
Substituting $g_{mt}$ from (A.3) back into the second and third equations in (A.2) gives two conditions describing savings and inculcation behavior:
$-\frac{1}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}}+\beta\frac{R_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$ (A.4)
@@
The two first-order conditions in (A.4) imply that for both savings and inculcation to occur, it must be the case that:
$\Psi_{t+1}=\frac{d_{t}R_{t+1}}{1+n}$ (A.5)
Multiplying both sides of (A.5) by $\eta_{t+1}$ then indicates that $g_{mt+1}$ in A.1 can be written as:
$g_{mt+1}=\eta_{t+1}\frac{d_{t}R_{t+1}}{(1+n)}$ (A.6)
Since $g_{et+1}=(1+n)g_{mt+1}$, we find that $g_{et+1}=\eta_{t+1}d_{t}R_{t+1}$. Using this in either part of (A.4) gives an expression describing how total savings, in effect, are distributed between asset accumulation $s_{mt}$ and inculcation $\eta_{t+1}$, which is reported in the text as equation (37):
$s_{mt}+\eta_{t+1}d_{t}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (A.7)
Expression (A.7) says that the typical middle-aged agent desires to save a fraction of current income for future savings. This savings is divided between actual savings, $s_{mt}$, and inculcative cultural investments, $\eta_{t+1}d_{t}$. Savings is also reduced according to the current discounted value of future income earned when elderly. The equilibrium condition (A.5) implies that if an agent is to engage in both asset accumulation and inculcation, the returns to each activity must be equal. Inserting the right-hand side of (A.7) into the right-hand side of (A.3) gives the equilibrium gift function in the text, equation (36):
$g^{*}_{mt}=\frac{\eta_{t}}{1+\beta}\left(y_{mt}+\frac{y_{et+1}}{R_{t+1}}\right)$ (A.8)
Comparing (A.1) and (A.8), we see that
$\Psi_{t}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (A.9)
Appendix B Proof of Proposition 5
The first part of the proposition concerning aggregate savings follows directly from Appendix A, and, in particular, Equation (A.7). To get the second part of the Proposition concerning returns on savings, iterate (A.5) back one period and substitute in the value of $\Psi_{t}$ so obtained into (A.9) to get the rate of return equalisation condition:
$R_{t}=\left(\frac{1+n}{1+\beta}\right)\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{d_{t-1}}$ (B.10)
Which (B.10) appears in the text as (38).
@@ Appendix C Proof of Proposition 6
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{y_{et+1}}{\delta y_{mt}}\frac{1}{R_{t+1}(1+\beta)}-\frac{s_{mt}}{\delta y_{mt}}$ (C.1)
Substitution for $R_{t+1}$, $y_{mt}$ and $y_{et+1}$ using (40) in (C.1), and also noting that the future capital stock depends on savings by middle-aged agents in the present, $K_{t+1}=s_{mt}N_{mt}$, gives:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et+1}k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{N_{mt}\delta\alpha A_{mt}k_{t}^{1-\alpha}}$ (C.2)
Multiplying the last term on the right by $L_{t+1}/L_{t+1}$, and substituting $N_{mt+1}=(1+n)N_{mt}$, $A_{mt+1}=(1+a)A_{mt}$, and $A_{et+1}=(1+a)A_{et}$ into (C.2), we have:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et}(1+a)k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{L_{t+1}}\frac{(1+a)(1+n)}{\frac{N_{mt+1}A_{mt+1}}{L_{t+1}}\delta\alpha k_{t}^{1-\alpha}}$ (C.3)
@@
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{k_{t+1}}{k_{t}^{1-\alpha}}\left(\frac{\tau_{e}(1+a)}{(1-\alpha)(1+\beta)\delta}+\frac{(1+n)(1+a)}{\alpha\delta(1-\gamma_{e})}\right)$ (C.4)
@@
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{A_{mt}k_{t}^{1-\alpha}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}+\frac{\frac{A_{et+1}k_{t+1}^{1-\alpha}}{R_{t+1}}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}\right)$ (C.5)
@@
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}\left(\frac{k_{t}}{k_{t-1}}\right)^{1-\alpha}+\frac{r_{e}(1+a)^{2}}{R_{t+1}\delta}\left(\frac{k_{t+1}}{k_{t-1}}\right)^{1-\alpha}\right)$ (C.6)
Putting in (C.6) the expression for asset returns, $R_{t}=(1-\alpha)k_{t}^{-\alpha}$, gives expression (42) in the text. One convenience of equation (C.6) is that it yields a simple steady-state relationship for asset returns:
$R=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}+\frac{(1+a)^{2}r_{e}}{\delta R}\right)$ (C.7)
@@
Appendix D Proof of Proposition 7
In this appendix, we derive the relative consumption ratio for both the case in which inculcation occurs and when it does not, starting with the latter case. In this case, elderly consumption is current income plus income earned from asset accumulation in the prior period:
$c_{et}=y_{et}+s_{mt-1}R_{t}$ (D.1)
@@ Appendix D Proof of Proposition 7
$s_{mt}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (D.3)
Plugging (D.3) into (D.2) gives:
$c_{mt}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (D.4)
One can also use (D.3) to deduce that elderly consumptio in (D.1) can be written as:
$c_{et}=\frac{\beta}{1+\beta}\left(y_{et}+R_{t}y_{mt-1}\right)$ (D.5)
@@
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\left(\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}\right)$ (D.6)
@@
$c_{mt}=y_{mt}-g_{mt}-s_{mt}-d_{t}\eta_{t+1}$ (D.8)
Elderly consumption is most easily found using (A.4), which allows us to conclude that:
$s_{mt-1}R_{t}+(1+n)g_{mt}+y_{et}=\beta R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (D.9)
Using (D.9) in (D.7) gives:
$c_{et}=R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (D.10)
Eliminating gifts in (D.8) using (36) gives:
$c_{mt}=(1-\eta)(y_{mt}-s_{mt}-d_{t}\eta_{t+1})$ (D.11)
For both (D.9) or (D.11), we can apply (37). Then, forming the relative consumption ratio gives:
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}$ (D.12)

### v2

@@ Appendix B Proof of Proposition 4
The solution to (35) starts by assuming an interior solution and then checking when this solution works. Assume that agents maintain the belief that gifts are a multiplicative function of the value of $\eta_{t}$:
$g_{mt+1}=\eta_{t+1}\Psi_{t+1}$ (B.1)
@@
where $\Psi_{t+1}$ is to be determined. Substituting for $g_{mt+1}$ using (B.1) into the middle-aged agent’s optimization problem (35) and differentiating gives first-order conditions:
$-\frac{1-\eta_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}}+\frac{\eta_{t}}{g_{mt}}$ $=0$ (B.2)
@@
Rearranged, the first expression in (B.2) implies that gifts are proportional to income:
$g_{mt}=\eta_{t}(y_{mt}-d_{t}\eta_{t+1}-s_{mt})$ (B.3)
Substituting $g_{mt}$ from (B.3) back into (B.2) gives two conditions describing savings and inculcation behavior:
$-\frac{1}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}}+\beta\frac{R_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$ (B.4)
@@
Equations (B.4) imply that for both savings and inculcation to occur, it must be that:
$\Psi_{t+1}=\frac{d_{t}R_{t+1}}{1+n}$ (B.5)
Multiplying both sides of (B.5) by $\eta_{t+1}$ gives $g_{mt+1}$ in B.1:
$g_{mt+1}=\eta_{t+1}\frac{d_{t}R_{t+1}}{(1+n)}$ (B.6)
Since $g_{et+1}=(1+n)g_{mt+1}$, we find that $g_{et+1}=\eta_{t+1}d_{t}R_{t+1}$. Using this in either part of (B.4) gives an expression describing how total savings are distributed between asset accumulation $s_{mt}$ and inculcation $\eta_{t+1}$, which is equation (37) in the text:
$s_{mt}+\eta_{t+1}d_{t}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (B.7)
Inserting the right-hand side of (B.7) into the right-hand side of (B.3) gives the equilibrium gift function in the text, equation (36):
$g^{*}_{mt}=\frac{\eta_{t}}{1+\beta}\left(y_{mt}+\frac{y_{et+1}}{R_{t+1}}\right)$ (B.8)
Comparing (B.1) and (B.8), we see that
$\Psi_{t}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (B.9)
@@ Appendix C Proof of Proposition 5
The first part of the proposition concerning aggregate savings follows directly from Appendix B, equation (B.7). To get the second part of the Proposition, iterate (B.5) back one period and substitute the value of $\Psi_{t}$ into (B.9) to get:
$R_{t}=\left(\frac{1+n}{1+\beta}\right)\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{d_{t-1}}$ (C.1)
Which (C.1) appears in the text as (38).
Appendix D Proof of Proposition 6
@@
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{y_{et+1}}{\delta y_{mt}}\frac{1}{R_{t+1}(1+\beta)}-\frac{s_{mt}}{\delta y_{mt}}$ (D.1)
Substitution for $R_{t+1}$, $y_{mt}$ and $y_{et+1}$ using (40) in (D.1), and also noting that the future capital stock depends on savings by middle-aged agents in the present, $K_{t+1}=s_{mt}N_{mt}$, gives:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et+1}k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{N_{mt}\delta\alpha A_{mt}k_{t}^{1-\alpha}}$ (D.2)
Multiplying the last term on the right by $L_{t+1}/L_{t+1}$, and substituting $N_{mt+1}=(1+n)N_{mt}$, $A_{mt+1}=(1+a)A_{mt}$, and $A_{et+1}=(1+a)A_{et}$ into (D.2), we have:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et}(1+a)k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{L_{t+1}}\frac{(1+a)(1+n)}{\frac{N_{mt+1}A_{mt+1}}{L_{t+1}}\delta\alpha k_{t}^{1-\alpha}}$ (D.3)
@@
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{k_{t+1}}{k_{t}^{1-\alpha}}\left(\frac{\tau_{e}(1+a)}{(1-\alpha)(1+\beta)\delta}+\frac{(1+n)(1+a)}{\alpha\delta(1-\gamma_{e})}\right)$ (D.4)
@@
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{A_{mt}k_{t}^{1-\alpha}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}+\frac{\frac{A_{et+1}k_{t+1}^{1-\alpha}}{R_{t+1}}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}\right)$ (D.5)
@@
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}\left(\frac{k_{t}}{k_{t-1}}\right)^{1-\alpha}+\frac{r_{e}(1+a)^{2}}{R_{t+1}\delta}\left(\frac{k_{t+1}}{k_{t-1}}\right)^{1-\alpha}\right)$ (D.6)
Putting in (D.6) the expression for asset returns, $R_{t}=(1-\alpha)k_{t}^{-\alpha}$, gives expression (42) in the text. One convenience of equation (D.6) is that it yields a simple steady-state relationship for asset returns:
$R=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}+\frac{(1+a)^{2}r_{e}}{\delta R}\right)$ (D.7)
@@
Appendix E Proof of Proposition 7
We derive the relative consumption ratio for both the case in which inculcation occurs and when it does not, starting with the latter case. In this case, elderly consumption is current income plus income earned from asset accumulation in the prior period:
$c_{et}=y_{et}+s_{mt-1}R_{t}$ (E.1)
@@ Appendix E Proof of Proposition 7
$s_{mt}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (E.3)
Plugging (E.3) into (E.2) gives:
$c_{mt}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (E.4)
One can also use (E.3) to deduce that elderly consumptio in (E.1) can be written as:
$c_{et}=\frac{\beta}{1+\beta}\left(y_{et}+R_{t}y_{mt-1}\right)$ (E.5)
@@
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\left(\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}\right)$ (E.6)
@@
$c_{mt}=y_{mt}-g_{mt}-s_{mt}-d_{t}\eta_{t+1}$ (E.8)
Elderly consumption is most easily found using (B.4), which allows us to conclude that:
$s_{mt-1}R_{t}+(1+n)g_{mt}+y_{et}=\beta R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (E.9)
Using (E.9) in (E.7) gives:
$c_{et}=R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (E.10)
Eliminating gifts in (E.8) using (36) gives:
$c_{mt}=(1-\eta)(y_{mt}-s_{mt}-d_{t}\eta_{t+1})$ (E.11)
For both (E.9) or (E.11), we can apply (37). Then, forming the relative consumption ratio gives:
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}$ (E.12)